OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

Trumpet, I can't figure it out. xD

OpenStudy (solomonzelman):

y + 3 = −5(x − 2) ---------> y=-5x+7 y − 2 = −5(x + 3) --------> y=-5x+13 y + 3 = −one fifth(x − 2) ---> y=1/5x-13/5 y − 2 = −one fifth(x + 3) ---> y=-1/5x+7/5

OpenStudy (anonymous):

instead of one fifth, you could also multiply both sides by 5 because you know.. well, for me its not uh good. isnt nice seeing a fraction form in an equation..

OpenStudy (anonymous):

@SolomonZelman right :)

OpenStudy (solomonzelman):

@BlveXO, find the slope of (−3, 2) and (2, 1).